# UniRAG: Universal Retrieval Augmentation for Multi-Modal Large Language Models

### v2

@@ 4.3 Configuration Details
Table 4: Caption generation with LLaVA on the Fashion200k using UniIR’s CLIP-SF and BLIP-FF as retrievers. Column $k\in\{0,1,5\}$ represents the number of few-shot examples for each experiment.
5 Evaluation Results and Analysis
Tables 36 present results for caption and image generation tasks on MSCOCO and Fashion200k datasets. Rows 1* from each table measure the effectiveness of CLIP-SF and BLIP-FF retrievers. They reflect how well the generator models perform if they return the image or caption from the first retrieved example as is. To ensure an equal number of images and captions for evaluation, we report retrievers’ performance using only their top retrieved image-text pair ($k=1$) for each query.
@@ 5 Evaluation Results and Analysis
The remaining rows in each table compare zero-shot ($k=0$) and few-shot ($k>0$) results of generator models using CLIP-SF (*a), BLIP-FF (*b), and random selection (*c) as retrievers for few-shot generations. The performance of CLIP-SF and BLIP-FF retrievers in RAG applications aligns with their standalone effectiveness (comparing rows *a and *b in each table). On MSCOCO, both retrievers continue to perform similarly, while BLIP-FF consistently outperforms CLIP-SF on Fashion200k.
@@ 5.1 Caption Generation
As reported in Table 3, augmenting relevant image-caption pairs ($k>0$) retrieved from the previous stage, boosts the effectiveness of MM-LLMs regardless of their baseline effectiveness, the number of in-context examples and the choice of retriever model. The baseline effectiveness ($k=0$) of LLaVA exceeds that of Gemini-Pro and GPT-4o models (comparing rows 2-4). Even though all models benefit from RA, this gap further widens when only a single retrieved example ($k=1$) is added to the prompts, making one-shot BLIP-FF + LLaVA the most effective setting and one-shot CLIP-SF + LLaVA a very close runner up. Further increasing the number of retrieved examples from one to five helps the Gemini-P and GPT-4o models. However, LLaVA with five retrieved examples is less effective than LLaVA with a single example.
Table 4 shows the caption generation results for the Fashion200k dataset. To reduce the cost of API calls, we only report results for the LLaVA model in this experiment. LLaVA performs poorly in zero-shot scenarios due to its limited knowledge of the domain-specific Fashion200k dataset. While randomly selected few-shot examples can help LLaVA grasp the task intent and the required level of detail for describing fashion products, the real improvement occurs when relevant examples are included in the prompts. Because many entities in the Fashion200k dataset are quite similar, the generator often repeats the in-context caption of a similar image for its query image, leading to effectiveness that is close to that of the baseline retriever models. Please see Appendix E for sample visualizations.
5.2 Image Generation
@@
Table 6: Image generation with LaVIT on the Fashion200k’s dataset using UniIR’s CLIP-SF and BLIP-FF as retrievers. Column $k\in\{0,1,5\}$ represents the number of few-shot examples for each experiment. For Inception Score (IS), its Standard Deviation (SD) is reported in parenthesis.
@@
Table 6 presents the image generation results for LaVIT on the Fashion200k dataset. As expected, FID and CLIP Score show trends similar to those in image captioning, with random and RAG examples incrementally enhancing the fidelity of generated images to the ground truth (lower FID) and improving the alignment of generated images with caption queries (higher CLIP Score). However, the Inception Score (IS) is highest in the zero-shot setting. This is due to the low diversity of the Fashion200k dataset; generating more faithful images in this context leads to lower diversity among the generated images, resulting in a lower IS. Another notable observation is that, with RAG examples, LaVIT achieves a CLIP Score higher than those of the retrievers. This suggests that the generated images align better with the caption queries than the images retrieved from the dataset. Please see Appendix E for sample visualizations.
5.2.1 Effect of Sampling
For each image in M-BEIR’s MSCOCO test set, we have sampled a single caption (Set 1 Captions in Table 7). To study the sensitivity to sampling, we repeated the sampling process one more time (Set 2 Captions in Table 7). The two sample sets have only 1030 captions in common but both of them cover all 5k distinct images in the dataset. We used CLIP-SF + LaVIT in this ablation study as the retriever-generator pipeline.
@@ 6 Conclusion
We introduced UniRAG as a model agnostic retrieval augmentation technique for Multi-Modal (MM) tasks with image and text modalities and evaluated it on image-to-text and text-to-image tasks. We utilized the LLaVA, Gemini-Pro, and GPT-4o models to generate captions for input images in zero-shot and few-shot (with RAG) settings. Similarly, we employed the LaVIT and Emu2-Gen models to generate images from input captions. In the RAG few-shot setting, we leveraged UniIR’s CLIP-SF and BLIP-FF models to retrieve relevant image-text pairs and included them as in-context examples. To further showcase the effectiveness of RAG few-shot examples, we also compared them against randomly selected few-shot examples.
Our experimental results on the MSCOCO dataset from M-BEIR indicated that all models, regardless of their zero-shot baseline effectiveness, improved after being exposed to in-context examples. However, the best performance was achieved only when relevant retrieved examples were included in the prompts, rather than random text-image pairs. In fact, for the LLaVA, GPT-4o, and Emu2-Gen models, adding five random examples decreased generation quality below their zero-shot baselines. In contrast, in the RAG few-shot setting, increasing the number of examples from one to five either improved generation quality (for Gemini-Pro and GPT-4o) or had no effect (for LaVIT and Emu2-Gen). For LLaVA, while adding more RAG in-context examples reduced its effectiveness, it still performed significantly better than both the zero-shot and random few-shot prompts. Our experiments on five MM-LLMs confirm that UniRAG with UniIR retrievers is a model-agnostic effective way to improve the generation quality of pre-trained MM-LLMs by retrieving relevant few-shot examples at inference time.
@@ 7 Limitations
While we open-sourced UniRAG, its usage is still bound by the licenses and the usage agreements of proprietary GPT-4o and Gemini-Pro models used for inference. Furthermore, one of our open-source models LaVIT, is licensed under the LaVIT Community License which requires specific license requests for commercial use.
This work uses English-only MSCOCO and Fashion200k datasets for evaluation and experiments with retriever and generator models that are primarily trained on English corpus. Thus, the effectiveness of our proposed method for non-English low-resource languages remains unexplored. Furthermore, relevance is the only criterion used while retrieving candidate pairs for in-context examples. Deployment of this method in real-life applications requires candidate rankings based on facts, potential harms, biases and other important factors for responsible AI, rather than solely relying on relevance.
Additionally, generating images of real people would require extra caution and extensive auditing since it has significant privacy and security implications. Finally, further research is required to investigate the generalizability of the retriever-guided generation to out-of-domain retrieval. For instance, employing an entity-centric evaluation dataset, which the retriever has not been trained on, could more effectively demonstrate the advantage of retrieval augmentation for multi-modal inference.
@@ Appendix B Proprietary Models Cost
Table 8: Model API pricing for per 1000 tokens. Above mentioned rates are in terms of USD. * - Gemini API call for image input presents cost per image.
@@ Appendix C Generation Visualization
Appendix D Reported Metrics
For image generation, we use FID Heusel et al. (2017) to measure the KL divergence between the feature vector distributions of generated and ground-truth images; a lower FID indicates greater similarity. Additionally, we calculate the CLIP Score Hessel et al. (2021), which measures cosine similarity between CLIP’s visual and text embeddings for a generated image and its caption query. Lastly, we include the Inception Score (IS) Salimans et al. (2016) and its Standard Deviation (SD) as another quality indicator for generated images. A higher IS indicates more diversity and confident classification of the generated images. However, IS provides a weaker signal compared to the other two metrics for image generation, as it is less accurate for small datasets.
@@ Appendix D Reported Metrics
Zeroshot CLIP-SF BLIP-FF Random
@@
Zeroshot CLIP-SF BLIP-FF Random
@@
Table 10: Sample image generation with LaVIT on Fashion200k dataset in zero-shot, and one-shot settings. The “Prompt” row shows the zero-shot caption query as well as retrieved image-caption pairs from CLIP-SF, BLIP-FF and random selection that are included in the prompt as in-context examples.
Appendix E Fashion200k Dataset Visualization
Tables 9 and 10 visualize sample caption and image generation for the Fashion200k dataset.

### v3

@@ 5 Evaluation Results and Analysis
Table 4: Caption generation with LLaVA on the Fashion200k dataset using UniIR’s CLIP-SF and BLIP-FF as retrievers. Column $k\in\{0,1,5\}$ represents the number of few-shot examples for each experiment.
Tables 3, 4, 5 and 6 present results for caption and image generation tasks on MSCOCO and Fashion200k datasets. Rows 1* from each table measure the effectiveness of CLIP-SF and BLIP-FF retrievers. They reflect how well the generator models perform if they return the image or caption from the first retrieved example as is. To ensure an equal number of images and captions for evaluation, we report retrievers’ effectiveness using only their top retrieved image-text pair ($k=1$) for each query.
@@
The remaining rows in each table compare zero-shot ($k=0$) and few-shot ($k>0$) results of generator models using CLIP-SF (*a), BLIP-FF (*b), and random selection (*c) as retrievers for few-shot generations. The behavior of CLIP-SF and BLIP-FF retrievers in RAG applications matches their standalone effectiveness (comparing rows *a and *b in each table). On MSCOCO, both retrievers continue to perform similarly, while BLIP-FF consistently outperforms CLIP-SF on Fashion200k.
@@ 5.1 Caption Generation
As reported in table 3, augmenting relevant image-caption pairs ($k>0$) retrieved from the previous stage, boosts the effectiveness of LVLMs regardless of their baseline effectiveness, the number of in-context examples and the choice of retriever model. The baseline effectiveness ($k=0$) of LLaVA exceeds that of Gemini-Pro and GPT-4o models (comparing rows 2-4). Even though all models benefit from RA, this gap further widens when only a single retrieved example ($k=1$) is added to the prompts, making one-shot BLIP-FF + LLaVA the most effective setting and one-shot CLIP-SF + LLaVA a very close runner up. Further increasing the number of retrieved examples from one to five helps the Gemini-P and GPT-4o models. However, LLaVA with five retrieved examples is less effective than LLaVA with a single example.
@@
Table 6: Image generation with LaVIT on the Fashion200k’s datasetg UniIR’s CLIP-SF and BLIP-FF as retrievers. Column $k\in\{0,1,5\}$ represents the number of few-shot examples for each experiment. For Inception Score (IS), its Standard Deviation (SD) is reported in parenthesis.
Table 4 shows the caption generation results for the Fashion200k dataset. To reduce the cost of API calls, we only report results for the LLaVA model in this experiment. LLaVA performs poorly in zero-shot scenarios due to its limited knowledge of the domain-specific Fashion200k dataset. While randomly selected few-shot examples can help LLaVA grasp the task intent and the required level of detail for describing fashion products, the real improvement occurs when relevant examples are included in the prompts. Because many entities in the Fashion200k dataset are quite similar, the generator often repeats the in-context caption of a similar image for its query image, leading to effectiveness that is close to that of the baseline retriever models. Please see Appendix D for sample visualizations.
5.2 Image Generation
@@
Table 6 presents the image generation results for LaVIT on the Fashion200k dataset. As expected, FID and CLIP Score show trends similar to those in Table 3, with random and RAG examples incrementally enhancing the fidelity of generated images to the ground-truth (lower FID) and improving the alignment of generated images with caption queries (higher CLIP Score). However, the Inception Score (IS) is highest in the zero-shot setting. This is due to the low diversity of the Fashion200k dataset; generating more faithful images in this context leads to lower diversity among the generated images, resulting in a lower IS. Another notable observation is that, with RAG examples, LaVIT achieves a CLIP Score higher than those of the retrievers. This suggests that the generated images align better with the caption queries than the images retrieved from the dataset. Please see Appendix D for sample visualizations.
5.3 Effect of Sampling
For each image in M-BEIR’s MSCOCO test set, we have sampled a single caption (Set 1 Captions in Table 7) as a query. To study the sensitivity to sampling, we repeated the sampling process one more time (Set 2 Captions in Table 7). The two sample sets have only 1030 captions in common but both of them cover all 5k distinct images in the dataset. We used CLIP-SF + LaVIT in this ablation study as the retriever-generator pipeline.
Baseline MSCOCO Fashion200k
CLIP-SF + Ground-truth 23.39 36.80
      LaVIT Retrieved 25.49 37.13
BLIP-FF + Ground-truth 23.23 36.68
      LaVIT Retrieved 25.17 36.34
Table 8: Comparison of FID measured in two modes: (a) generated images vs. retrieved images and (b) generated images vs. ground-truth. The images are generated by LaVIT using one-shot examples retrieved by CLIP-SF and BLIP-FF retrievers.
5.4 Datasets Comparison
As shown in Tables 4 and 6, smaller models perform poorly on the Fashion200k dataset in both tasks under zero-shot settings, indicating their limited implicit knowledge of the fashion domain. However, their strong reliance on in-context examples and significant improvement under few-shot RAG settings demonstrate UniRAG’s effectiveness in bridging this knowledge gap.
In contrast, for the MSCOCO dataset (Tables 3 and 5), where smaller models have more inherent knowledge of common entities, the zero-shot effectiveness is greater and the impact of few-shot RAG examples is less pronounced. While few-shot RAG examples provide greater benefits for the Fashion200k dataset, this does not imply that the model merely replicates retrieved images in the absence of pre-trained knowledge. As shown in Table 8, the Fashion200k dataset exhibits a significantly higher FID when comparing generated images to the top-one retrieved image, indicating that the model does not simply copy but instead generates distinct outputs.
@@ 6 Conclusion
We introduced UniRAG as a model agnostic retrieval augmentation technique for Vision-Language (VL) tasks and evaluated it on image-to-text and text-to-image tasks. We utilized the LLaVA, Gemini-Pro, and GPT-4o models to generate captions for input images in zero-shot and few-shot (with RAG) settings. Similarly, we employed the LaVIT and Emu2-Gen models to generate images from input captions. In the RAG few-shot setting, we leveraged UniIR’s CLIP-SF and BLIP-FF models to retrieve relevant image-text pairs and included them as in-context examples. To further showcase the effectiveness of RAG few-shot examples, we also compared them against randomly selected few-shot examples.
Our experimental results on the MSCOCO dataset from M-BEIR indicated that all models, regardless of their zero-shot baseline effectiveness, improved after being exposed to in-context examples. However, the best results were achieved only when relevant retrieved examples were included in the prompts, rather than random text-image pairs. In fact, for the LLaVA, GPT-4o, and Emu2-Gen models, adding five random examples decreased generation quality below their zero-shot baselines. In contrast, in the RAG few-shot setting, increasing the number of examples from one to five either improved generation quality (for Gemini-Pro and GPT-4o) or had no effect (for LaVIT and Emu2-Gen). For LLaVA, while adding more RAG in-context examples reduced its effectiveness, it still performed significantly better than both the zero-shot and random few-shot prompts. Our experiments on five LVLMs confirmed that UniRAG with UniIR retrievers is a model-agnostic effective way to improve the generation quality of pre-trained LVLMs by retrieving relevant few-shot examples at inference time.
@@ 7 Limitations
While we have open-sourced UniRAG, its usage is still bound by the licenses and the usage agreements of proprietary GPT-4o and Gemini-Pro models used for inference. Furthermore, one of our open-source models LaVIT, is licensed under the LaVIT Community License which requires specific license requests for commercial use.
The primary goal of UniRAG is to enhance the model’s knowledge of specific subject entities, ensuring that retrieved information directly influences the generated output—a common characteristic of all RAG applications. While retrieving from a diverse database and incorporating multiple few-shot examples can broaden this influence, this approach may be less suitable for applications prioritizing creativity or originality over generation accuracy. Additionally, relevance is the only criterion used while retrieving candidate pairs for in-context examples. Deploying this method in real-world applications requires ranking candidates based on facts, potential harms, biases and other important factors for responsible AI, rather than solely relying on relevance. The added latency and computation cost of retrieval and inference with larger prompts must also be considered for time- or cost-sensitive applications.
This work uses the English-only MSCOCO and Fashion200k datasets for evaluation and experiments with retriever and generator models that are primarily trained on English corpus. Thus, the effectiveness of our proposed method for non-English low-resource languages remains unexplored. Moreover, generating images of real people necessitates extra caution and extensive auditing due to significant privacy and security implications. Finally, further research is required to assess the generalizability of the retriever-guided generation to out-of-domain retrieval. For instance, employing an entity-centric evaluation dataset, which the retriever has not been trained on, could better demonstrate the advantage of retrieval augmentation for multi-modal inference.
Acknowledgments
This research was supported in part by the Natural Sciences and Engineering Research Council (NSERC) of Canada. We also thank Microsoft and Google for providing access to OpenAI LLMs via Azure and Gemini via VertexAI, respectively. Additionally, we appreciate Cong Wei for answering several questions about UniIR and Xueguang Ma for providing valuable feedback.
@@ Appendix A Prompt Templates
Table 9: Model API pricing per 1,000 tokens is listed in USD. (*) The cost for a Gemini API call with image input is presented per image.
Appendix B Proprietary Models Cost
@@ Appendix C Generation Visualization
Zero-shot CLIP-SF BLIP-FF Random
@@
Zero-shot CLIP-SF BLIP-FF Random
@@
Table 11: Sample image generation with LaVIT on the Fashion200k dataset in zero-shot and one-shot settings. The “Prompt” row shows the zero-shot caption query as well as retrieved image-caption pairs from CLIP-SF, BLIP-FF and random selection that are included in the prompt as in-context examples.
Appendix D Reported Metrics
For image generation, we use FID Heusel et al. (2017) to measure the KL divergence between the feature vector distributions of generated and ground-truth images; a lower FID indicates greater similarity. Additionally, we calculate the CLIP Score Hessel et al. (2021), which measures cosine similarity between CLIP’s visual and text embeddings for a generated image and its caption query. Lastly, we include the Inception Score (IS) Salimans et al. (2016) and its Standard Deviation (SD) as another quality indicator for generated images. A higher IS indicates more diversity and confident classification of the generated images. However, IS provides a weaker signal compared to the other two metrics for image generation, as it is less accurate for small datasets. onFashion200k Dataset Visualization Tables 10 and 11 visualize sample caption and image generation for the Fashion200k dataset.